OpenStudy (anonymous):

simplify √18x^3/√3xy

OpenStudy (pfenn1):

\[\frac{\sqrt{18x^3}}{\sqrt{3xy}}=\sqrt{\frac{2(3)3x^3}{3xy}}\]Can you simplify this from here?

OpenStudy (anonymous):

dont mean to be an idiot but can you please just answer it lol

OpenStudy (pfenn1):

You can do it! You are not an idiot. Would it help if I put it this way?\[\sqrt{\frac{2(3)3x^3}{3xy}}=\sqrt{\left( \frac6y \right)\ \left( \frac33 \right)\left( \frac{x^3}{x} \right)}\]